Suppose that a trainee soldier shoots a target in an independent manner. If the probability that the



target is hit on any one shot is 0.8, what is the probability that the target would be hit



i. on the sixth attempt



ii. in fewer than 5 shots



iii. in even number of shots

1
Expert's answer
2021-12-07T10:13:08-0500

The Geometric distribution is

"P(X=x)=p(x)= q^{x-1}p,x= 1, 2, 3, ..."

Given "p=0.8, q=1-p=0.2"


i.


"P(X=6)=(0.2)^{6-1}(0.8)=0.000256"

ii.


"P(X<5)=P(X=1)+P(X=2)+P(X=3)"

"+P(X=4)=(0.2)^{1-1}(0.8)+(0.2)^{2-1}(0.8)"

"+(0.2)^{3-1}(0.8)+(0.2)^{4-1}(0.8)=0.9984"

iii.

"P(X\ is \ even\ number)=P(X=2)+P(X=4)"

"+P(X=6)+...=0.8(0.2)^{2-1}+0.8(0.2)^{4-1}"

"+0.8(0.2)^{6-1}=0.8(0.2+(0.2)^3+(0.2)^5+...)"

"=0.8(0.2)(\dfrac{1}{1-(0.2)^2})=\dfrac{1}{6}"